# The Mean of Multi-Object Trajectories

### v1

@@ I-D Proof of Proposition 3
[…]
Hence, for each $k$, $x_{k}^{*}(\gamma)=\arg\min_{x_{k}\in\mathbb{X}}\gamma_{k}\psi_{k}^{(r)}(\gamma,x_{k})$. If $\gamma_{k}\!=\!0$ or $\ \mathbf{1}_{\mathcal{D}_{\!v^{(n)}\!}}^{(k)}=\!0$, for all $n=\!1:N$, then any value of $x_{k}^{*}(\gamma)\!\in\!\mathbb{X}$ yields the same cost. Thus, $x_{k}^{*}(\gamma)\!=\!\boldsymbol{0}\in\mathbb{X}$ is an optimal solution. If $\gamma_{k}\!=\!1$, then $\gamma_{k}\psi_{k}^{(r)}(\gamma,\cdot)=\psi_{k}^{(r)}(\gamma,\cdot)$, and
[…]
I-E Proof of Proposition 4
If $L\triangleq\sum_{n=1}^{N}|\boldsymbol{Y}^{(n)}|=0$, an optimal solution (of Problem (17)-(LABEL:eq:ospa-based-traj-ori)) is $\hat{\boldsymbol{Y}}=\emptyset$, yielding zero cost.
@@ I-F Proof of Proposition 5
  • $\varepsilon_{\ell}(\boldsymbol{X})=1\Rightarrow\varepsilon_{\tilde{\pi}(\ell)}(\boldsymbol{Y}^{(n)\!})=1$ , if $\left\|\varepsilon(\boldsymbol{X})\right\|_{1}\leq||\varepsilon(\boldsymbol{Y}^{(n)\!})||_{1}$ ; or
  • $\varepsilon_{\tilde{\pi}(\ell)}(\boldsymbol{Y}^{(n)\!})=1\Rightarrow\varepsilon_{\ell}(\boldsymbol{X})=1$ , if $\left\|\varepsilon(\boldsymbol{X})\right\|_{1}>||\varepsilon(\boldsymbol{Y}^{(n)\!})||_{1}$
@@
$\tilde{d}_{\mathbb{T}}^{(c)}\!(\boldsymbol{X},\boldsymbol{Y}|\ell,\tilde{\pi})=\begin{cases}d_{\mathbb{T}}^{(c)}\!(\boldsymbol{X}_{\!\ell},\boldsymbol{Y}_{\!\!\tilde{\pi}(\ell)}),&\text{$\ell$$\leq|\boldsymbol{X}|$,$\tilde{\pi}$($\ell$)$\leq$|$\boldsymbol{Y}$|}\\ 0,&otherwise\end{cases}.$
@@ I-G Proof of Proposition 6
For each $\ell$, $\tau_{\ell}^{*}(\eta,\omega)=\arg\min_{\tau_{\ell}\in\mathbb{T}}\eta_{\ell}\phi_{\omega,\ell}^{(r)}(\eta,\tau_{\ell})$. If $\eta_{\ell}=0$ or $\xi_{\omega(n,\ell)}^{(n)}=0$ for all $n=\!1\!:\!N$, any value of $\tau_{\ell}^{*}(\eta,\omega)$ yields the same cost. Hence, $\tau_{\ell}^{*}(\eta,\omega)\!=\!\boldsymbol{0}$ is an optimal solution. If $\eta_{\ell}=1$, then $\eta_{\ell}\phi_{\omega,\ell}^{(r)}(\eta,\tau_{\ell})=\phi_{\omega,\ell}^{(r)}(\eta,\tau_{\ell})$, and
@@
$=\arg\min_{\tau_{\ell}\in\mathbb{T}}\sum_{n=1}^{N}\xi_{\omega(n,\ell)}^{(n)}\frac{d_{\mathbb{T}}^{(c)}(\tau_{\ell},\chi_{\omega(n,\ell)}^{(n)})^{r}}{\kappa(|\!|\eta|\!|_{1},|\!|\xi^{(n)}|\!|_{1})}.$
@@ II-A Convergence Properties
Case 2: Suppose $\left\|\eta_{1:L}\right\|_{1}<\left\|\xi^{(n)}\right\|_{1}$. Changing $\eta_{\ell}$ from 1 to 0 to reach an intermediate state $\grave{\eta}_{1:L}$ will satisfy $\omega^{(n)}\in\Omega^{(n)}(\grave{\eta}_{1:L})$, allowing a transition from $(\eta_{1:L},\omega^{(1:N)})$ to $(\grave{\eta}_{1:L},\omega^{(1:N)})$. If $\xi_{\ell}^{(n)}=1$, changing $\eta_{\ell}$ from 0 to 1 to reach $\grave{\eta}_{1:L}$ will also satisfy $\omega^{(n)}\in\Omega^{(n)}(\grave{\eta}_{1:L})$. However, if $\xi_{\ell}^{(n)}=0$, to change $\eta_{\ell}$ from 0 to 1 we first need to change $\omega^{(n)}$ to $\grave{\omega}^{(n)}$ as follows. For some $\grave{\ell}$ where $\eta_{\grave{\ell}}=0$ and $\xi_{\grave{\ell}}^{(n)}=1$, we select some $\grave{\omega}^{(n)}\in\Omega^{(n)}(\eta_{1:L})$ such that: $\grave{\omega}_{\ell}^{(n)}=\omega_{\grave{\ell}}^{(n)}$; $\grave{\omega}_{\grave{\ell}}^{(n)}=\omega_{\ell}^{(n)}$; and $\grave{\omega}_{\check{\ell}}^{(n)}=\omega_{\check{\ell}}^{(n)}$ for $\check{\ell}\neq\ell,\grave{\ell}$. Since $\left\|\eta_{1:L}\right\|_{1}<\left\|\xi^{(n)}\right\|_{1}$, there always exists a $\grave{\ell}$ that satisfies our requirement. Since $\grave{\omega}^{(n)}\in\Omega^{(n)}(\grave{\eta}_{1:L})$, it allows us to move from $(\eta_{1:L},\omega^{(1:N)})$ to $(\eta_{1:L},\grave{\omega}^{(1:N)})$ to $(\acute{\eta}_{1:L},\grave{\omega}^{(1:N)})$ and then $(\acute{\eta}_{1:L},\acute{\omega}^{(1:N)})$.
Case 3: Suppose $\left\|\eta_{1:L}\right\|_{1}>\left\|\xi^{(n)}\right\|_{1}$. Changing $\eta_{\ell}$ from 0 to 1 to reach an intermediate state $\grave{\eta}_{1:L}$ will satisfy $\omega^{(n)}\in\Omega^{(n)}(\grave{\eta}_{1:L})$, allowing a transition from $(\eta_{1:L},\omega^{(1:N)})$ to $(\grave{\eta}_{1:L},\omega^{(1:N)})$. If $\xi_{\ell}^{(n)}=0$, changing $\eta_{\ell}$ from 1 to 0 to reach $\grave{\eta}_{1:L}$ will also satisfy $\omega^{(n)}\in\Omega^{(n)}(\grave{\eta}_{1:L})$. However, if $\xi_{\ell}^{(n)}=1$, to change $\eta_{\ell}$ from 1 to 0 we first need to change $\omega^{(n)}$ to $\grave{\omega}^{(n)}$ as follows. For some $\grave{\ell}$ where $\eta_{\grave{\ell}}=1$ and $\xi_{\grave{\ell}}^{(n)}=0$, we select some $\grave{\omega}^{(n)}\in\Omega^{(n)}(\eta_{1:L})$ such that: $\grave{\omega}_{\ell}^{(n)}=\omega_{\grave{\ell}}^{(n)}$; $\grave{\omega}_{\grave{\ell}}^{(n)}=\omega_{\ell}^{(n)}$; and $\grave{\omega}_{\check{\ell}}^{(n)}=\omega_{\check{\ell}}^{(n)}$ for $\check{\ell}\neq\ell,\grave{\ell}$. Since $\left\|\eta_{1:L}\right\|_{1}>\left\|\xi^{(n)}\right\|_{1}$, there always exists a $\grave{\ell}$ that satisfies our requirement. Since $\grave{\omega}^{(n)}\in\Omega^{(n)}(\grave{\eta}_{1:L})$, it allows us to move from $(\eta_{1:L},\omega^{(1:N)})$ to $(\eta_{1:L},\grave{\omega}^{(1:N)})$ to $(\acute{\eta}_{1:L},\grave{\omega}^{(1:N)})$ and then $(\acute{\eta}_{1:L},\acute{\omega}^{(1:N)})$.
This implies the Markov chain is irreducible and regular, hence the convergence to the target distribution $\rho(\eta_{1:L},\omega^{(1:N)})$ at an exponential rate, following from Lemma 4.3.4 of [1]. $\blacksquare$
@@ II-B1 Rearranging Non-Existence Assignment
For some $\eta_{1:L}$, $\bar{L}=\sum_{\ell=1}^{L}(1-\eta_{\ell})$, we define the mapping ¯$\bar{\mathcal{E}}:\Pi_{L}\rightarrow\mathcal{F}_{\bar{L}}(\{1:L\})$ (where $\mathcal{F}_{\bar{L}}(\{1:L\})$ is all finite subsets of $\{1:L\}$ with $\bar{L}$ elements) such that ¯$\bar{\mathcal{E}}(\omega)=\{\omega_{\ell}|\eta_{\ell}=0,\ell=1:L\}$. For some $\omega\in\Pi_{L}$, we define the mapping $\mathcal{E}_{\omega}:\mathcal{F}_{\bar{L}}(\{1:L\})\rightarrow\Pi_{L}$ such that if $\mathcal{E}_{\omega}(\{\bar{\omega}_{1:\bar{L}}\})=\grave{\omega}$ then
@@
We note that if $\bar{\omega}\in\Pi_{\bar{\mathcal{E}}(\omega)}$ is a permutation of ¯$\bar{\mathcal{E}}(\omega)$ and $\omega\in\Omega^{(n)}(\eta_{1:L})$ then $\mathcal{E}_{\omega}(\bar{\omega})\in\Omega^{(n)}(\eta_{1:L})$. Thus, we can update $\omega^{(n)}\in\Omega^{(n)}(\eta_{1:L})$ to $\grave{\omega}^{(n)}\in\Omega^{(n)}(\eta_{1:L})$ using some $\bar{\omega}^{(n)}\in\Pi_{\bar{\mathcal{E}}(\omega^{(n)})}$ such that $\acute{\omega}_{\ell}^{(n)}=\mathcal{E}_{\omega^{(n)}}(\bar{\omega}^{(n)})$. Noting that $S^{(r)}(\eta_{1:L},\omega^{(1:N)})=S^{(r)}(\eta_{1:L},\grave{\omega}^{(1:N)})$. To achieve low pair-wise distances among $\chi_{\grave{\omega}_{\ell}^{(1)}}^{(n)}$, …, $\chi_{\grave{\omega}_{\ell}^{(N)}}^{(n)}$, we sample $\bar{\omega}^{(1:N)}$ from the distribution

### v2

@@ I-E Proof of Proposition 3
$U^{(r)}(\gamma,x^{*}(\gamma))$ $\!\!\!\!\!=\!\!\!\!\!\!$ $\min_{x\in\mathbb{X}^{|\mathbb{K}|}}{\displaystyle\sum_{k\in\mathbb{K}}}\gamma_{k}\psi_{k}^{(r)}(\gamma,x_{k})+\left(1-\gamma_{k}\right)\bar{\psi}_{k}^{(r)}(\gamma)$
$\!\!\!\!\!\!=\!\!\!\!\!\!$ ${\displaystyle\sum_{k\in\mathbb{K}}}\min_{x_{k}\in\mathbb{X}}\gamma_{k}\psi_{k}^{(r)}(\gamma,x_{k})\!+\!{\displaystyle\sum_{k\in\mathbb{K}}}\!\left(1\!-\!\gamma_{k}\right)\!\bar{\psi}_{k}^{(r)}(\gamma).$
@@
$\!\!\!\!\!\!x_{k}^{*}(\gamma)$ $\!\!\!=\!\!\!$ $\arg\min_{x_{k}\in\mathbb{X}}\psi_{k}^{(r)}(\gamma,x_{k})$
$\!\!\!=\!\!\!$ $\arg\min_{x_{k}\in\mathbb{X}}{\displaystyle\sum_{n=1}^{N}}\frac{\mathbf{1}_{\mathcal{D}_{\!v^{(n)}\!}}^{(k)}[d_{\mathbb{X}}^{(c)\!}(x_{k},v^{(n)\!}(k))]^{r}}{|\mathcal{D}_{\gamma}\cup\mathcal{D}_{v^{(n)}}|}.$
I-F Proof of Proposition 4
If $L\triangleq\sum_{n=1}^{N}|\boldsymbol{Y}^{(n)}|=0$, an optimal solution (of Problem (17)-(LABEL:eq:ospa-based-traj-ori)) is $\hat{\boldsymbol{Y}}$$\emptyset$, yielding zero cost.
@@ I-G Proof of Proposition 5
[…]
@@
$\tilde{d}_{\mathbb{T}}^{(c)}\!(\boldsymbol{X},\boldsymbol{Y}|\ell,\tilde{\pi})=\begin{cases}d_{\mathbb{T}}^{(c)}\!(\boldsymbol{X}_{\!\ell},\boldsymbol{Y}_{\!\!\tilde{\pi}(\ell)}),&\text{$\ell$$\leq|\boldsymbol{X}|$,$\tilde{\pi}$($\ell$)$\leq$|$\boldsymbol{Y}$|}\\ 0,&\textrm{otherwise}\end{cases}.$
@@ I-H Proof of Proposition 6
For each $\ell$, $\tau_{\ell}^{*}(\eta,\omega)=\arg\min_{\tau_{\ell}\in\mathbb{T}}\eta_{\ell}\phi_{\omega,\ell}^{(r)}(\eta,\tau_{\ell})$. If $\eta_{\ell}=0$ or $\xi_{\omega(n,\ell)}^{(n)}=0$ for all n=1:Nn=\!1\!:\!N, any value of $\tau_{\ell}^{*}(\eta,\omega)$ yields the same cost. Hence, $\tau_{\ell}^{*}(\eta,\omega)\!=\!\boldsymbol{0}$ is an optimal solution. If $\eta_{\ell}=1$, then $\eta_{\ell}\phi_{\omega,\ell}^{(r)}(\eta,\tau_{\ell})=\phi_{\omega,\ell}^{(r)}(\eta,\tau_{\ell})$, and
@@
[…]
@@ II-A Convergence Properties
[…]
This implies the Markov chain is irreducible and regular, hence the convergence to the target distribution $\rho(\eta_{1:L},\omega^{(1:N)})$ at an exponential rate, following from Lemma 4.3.4 of [gallager2013stochastic]. $\blacksquare$
@@ II-B1 Rearranging Non-Existence Assignment
For some $\eta_{1:L}$, $\bar{L}=\sum_{\ell=1}^{L}(1-\eta_{\ell})$, we define the mapping $\bar{\mathcal{E}}:\Pi_{L}\rightarrow\mathcal{F}_{\bar{L}}(\{1:L\})$ (where $\mathcal{F}_{\bar{L}}(\{1:L\})$ is all finite subsets of $\{1:L\}$ with $\bar{L}$ elements) such that $\bar{\mathcal{E}}(\omega)=\{\omega_{\ell}|\eta_{\ell}=0,\ell=1:L\}$. For some $\omega\in\Pi_{L}$, we define the mapping $\mathcal{E}_{\omega}:\mathcal{F}_{\bar{L}}(\{1:L\})\rightarrow\Pi_{L}$ such that if $\mathcal{E}_{\omega}(\{\bar{\omega}_{1:\bar{L}}\})=\grave{\omega}$ then
@@
We note that if $\bar{\omega}\in\Pi_{\bar{\mathcal{E}}(\omega)}$ is a permutation of $\bar{\mathcal{E}}(\omega)$ and $\omega\in\Omega^{(n)}(\eta_{1:L})$ then $\mathcal{E}_{\omega}(\bar{\omega})\in\Omega^{(n)}(\eta_{1:L})$. Thus, we can update $\omega^{(n)}\in\Omega^{(n)}(\eta_{1:L})$ to $\grave{\omega}^{(n)}\in\Omega^{(n)}(\eta_{1:L})$ using some $\bar{\omega}^{(n)}\in\Pi_{\bar{\mathcal{E}}(\omega^{(n)})}$ such that $\acute{\omega}_{\ell}^{(n)}=\mathcal{E}_{\omega^{(n)}}(\bar{\omega}^{(n)})$. Noting that $S^{(r)}(\eta_{1:L},\omega^{(1:N)})=S^{(r)}(\eta_{1:L},\grave{\omega}^{(1:N)})$. To achieve low pair-wise distances among $\chi_{\grave{\omega}_{\ell}^{(1)}}^{(n)}$, …, $\chi_{\grave{\omega}_{\ell}^{(N)}}^{(n)}$, we sample $\bar{\omega}^{(1:N)}$ from the distribution